# UniT: Unified Tactile Representation for Robot Learning

### v1

@@ Abstract
UniT is a novel approach to tactile representation learning, using VQVAE to learn a compact latent space and serve as the tactile representation. It uses tactile images obtained from a single simple object to train the representation with transferability and generalizability. This tactile representation can be zero-shot transferred to various downstream tasks, including perception tasks and manipulation policy learning. Our benchmarking on an in-hand 3D pose estimation task shows that UniT outperforms existing visual and tactile representation learning methods. Additionally, UniT’s effectiveness in policy learning is demonstrated across three real-world tasks involving diverse manipulated objects and complex robot-object-environment interactions. Through extensive experimentation, UniT is shown to be a simple-to-train, plug-and-play, yet widely effective method for tactile representation learning. For more details, please refer to our open-source repository [^1] https://github.com/ZhengtongXu/UniT and the project website [^2] https://zhengtongxu.github.io/unifiedtactile.github.io/.
@@ I Introduction
Imitation learning has shown promising results equipping robots in the domain of manipulation, with the ability to master complex, high-precision, dexterous skills through human demonstrations [1, 2, 3, 4]. However, the predominant focus of current research is on leveraging image or point cloud inputs to perceive scene information. During manipulation, robots encounter varied force interactions from both the environments and the manipulated objects. The sole dependence on images [1, 2, 3] or point clouds [4, 5] could obscure critical details about the objects’ in-hand states and the dynamics of force interactions, which may be crucial for effective manipulation. Despite the current focus on visual information in robot learning research, humans routinely use visual and tactile feedback in manipulation tasks. Therefore, exploring the integration of visual and tactile modalities in imitation learning could potentially enhance the robots’ performance in manipulation tasks.
Vision-based tactile sensors such as GelSight [6] offer high-resolution feedback that captures extensive tactile information. In tactile sensing, GelSight has demonstrated the ability to perceive the geometry of the grasped object [7], the position and orientation of in-hand objects [8, 9], the normal and shear force fields [10], and even the physical properties of objects [11, 12, 13]. Moreover, recent studies incorporating vision-based tactile sensors into imitation learning have demonstrated their considerable potential and improved performance [14, 15, 16]. Consequently, vision-based tactile sensors are highly effective for use in imitation learning for manipulation tasks.
Although incorporating visual-tactile feedback as policy input is a logical approach, it raises Question 1: Should we treat vision-based tactile feedback as typical visual input in robot learning? Treating it as such would imply that a system with tactile sensors is akin to one equipped with additional cameras, which could be processed by existing imitation learning frameworks.
However, does this approach truly capture the unique information of tactile feedback? Previous research indicates that the tactile images produced by GelSight, owing to their unique sensing principles, enable trained neural networks to exhibit strong generalizability. For example, a gradient estimation model trained solely on images from a small ball contacting the sensor can be generalized across objects with various shapes and textures [7, 17]. Similarly, a tactile-reactive grasping policy trained on standardized texture-less blocks shows generalizability to diverse objects with different shapes and textures [11]. Therefore, the generalizability demonstrated by GelSight highlights the distinctiveness of tactile images compared to standard visual images.
However, the tactile depth reconstruction and tactile reactive grasping models just discussed have only demonstrated generalizability within their respective specific tasks. Then, this leads to Question 2: Can we use a single simple object to learn a unified tactile representation that 1) possesses transferability and generalizability, and 2) incorporates as much of the rich information present in tactile images as possible? This representation would be advantageous in being applied in a zero-shot manner across a variety of downstream tasks involving different objects.
To address Question 1 and Question 2, we propose UniT, unified tactile representation for robot learning, as shown in Fig. LABEL:fig:first_page. Our contributions are summarized as follows:
@@
1. Easy to Train yet Broadly Applicable: We introduce UniT, a tactile representation learning method for GelSight sensors. UniT uses VQVAE to learn a compact latent space and serve as the tactile representation. It can utilize data on a single simple object to learn a unified tactile representation that can be generalized to objects of different sizes and shapes. Through experiments in image reconstruction involving diverse objects, we demonstrate that the representation learned by UniT from a single simple object can capture information of unseen objects on contact configurations, object shapes, and dynamic marker motions induced by applied forces. This kind of transferability and generalizability makes UniT both straightforward to train and broadly applicable.
@@
2. Deploy for Tactile Perception: The tactile encoder trained through UniT can be seamlessly transferred to downstream tactile perception tasks. Furthermore, due to UniT’s lenient data requirements during training, representations learned even from the simplest objects, such as a small ball, can effectively facilitate tactile perception of everyday objects. In experiments involving the task of estimating the in-hand 3D pose of the USB plug, UniT outperforms training of a ResNet [18] from scratch, existing visual representation learning methods BYOL [19] and MAE [20], and the state-of-the-art tactile representation learning method, T3 [21].
@@ II-A Tactile-involved Imitation Learning
Existing tactile-involved imitation learning can be broadly categorized into three types: those utilizing low-dimensional tactile feedback, such as tactile arrays or force-torque feedback; those employing audio feedback; and those utilizing high-dimensional, vision-based tactile feedback.
Research that incorporates tactile arrays [22, 23], force-torque feedback [24], and in-hand audio feedback [25] into imitation learning policies have shown strong performance in tasks with complex interactions between environments, objects, and robots. However, tactile interactions are inherently complex and these modalities often do not capture the full spectrum of tactile information. For example, tactile arrays may struggle to accurately perceive an object’s geometry, in-hand position, and orientation, while force-torque and in-hand audio feedback may not effectively capture dynamic force distributions on the hand/finger.
Additionally, vision-based tactile sensors such as GelSight [6] offer high-resolution feedback that captures extensive tactile information. Recent studies incorporating GelSight into imitation learning have demonstrated its considerable potential and enhanced performance [14, 15, 16].
@@ II-B Representation Learning in Imitation Learning and Tactile Sensing
Recent studies demonstrate that representation learning significantly enhances performance in imitation learning [1, 26, 27]. Encoders that are pretrained with representation learning are particularly adept at extracting information from visual observations.
In the field of tactile sensing, a variety of representation learning frameworks have been introduced, including the use of MAE [28, 29] and CNN [30]. Furthermore, the work in [21] introduces a framework for tactile representation learning that scales across multi-sensors and multi-tasks. The work in [31] learns multimodal tactile representation by aligning tactile embeddings to pretrained image embeddings associated with a variety of other modalities.
In this paper, we demonstrate that using tactile images from single simple object can be used to train a tactile representation with transferability and generalizability. This ability to “train simply but work broadly” is not possessed by existing tactile representation learning frameworks.
III Method
In this section, we detail our UniT pipeline and the integration of the pretrained encoders into downstream perception and policy learning tasks.
@@ III Method
III-A VQGAN Autoencoder
The work in [32] introduces VQGAN for high-resolution image synthesis. VQGAN integrates a vector quantized variational autoencoder (VQVAE) [33] with a patch-based discriminator. Subsequently, the work in [34] introduces latent diffusion model, which leverages the VQGAN autoencoder to compress images into a latent space where diffusion and denoising processes can occur. The efficacy of the latent diffusion model is largely due to the robust image compression capabilities of VQGAN, which not only facilitates high-quality image reconstruction but also avoids high degree of variance in the learned latent space through VQ regularization.
Although VQGAN demonstrates impressive performance in generative model and image generation areas, its role as an autoencoder is primarily to enable high-quality image compression and reconstruction. This function differs from the objectives of representation learning, which focus on capturing representations that distill the essential features of the original images (although reconstruction can be used to validate the effectiveness of representation learning). Such representations aim to generalize across a variety of downstream tasks, leveraging the unified features to perform diverse functions effectively. Therefore, for visual images, VQGAN may not necessarily be an ideal framework for representation learning.
Tactile images are derived from an imaging principle that combines RGB-tricolor light scattering with gel deformation. These images, produced by different objects, show different features. However, the color distribution of tactile images is significantly more compact than that of visual images. This compactness arises because tactile images are generated using the same imaging principles, which disregard the color of the object and background, focusing instead only on tactile features such as contact configuration, geometry, and force distribution. We believe that the compactness of tactile image distributions makes it possible to adequately represent these images in a more compact representation. Consequently, we utilize VQGAN for learning tactile representations.
@@ III-A VQGAN Autoencoder
As depicted in Fig. 1, the representation learning pipeline employs a VQGAN architecture with the quantization layer absorbed into the decoder. This pipeline comprises a CNN encoder ${E}$, a CNN decoder (incorporating the quantization layer) ${G}$, and a patch-based discriminator ${D}$. The entire framework is trained in a self-supervised manner. Denote the tactile image input as $x\in\mathbb{R}^{{H}\times{W}\times{3}}$, where the encoder ${E}$ maps the image to a tactile representation $z={E}(x)\in\mathbb{R}^{{h}\times{w}\times{c}}$. The decoder reconstructs $z$ to an image $\hat{x}$.
@@
The use of a discriminator is justified because recent studies [35, 36] have shown that adversarial loss can enhance the performance of representation learning. The presence of VQ regularization enables the learning of a compact tactile representation with reduced variance, which can improve the performance of representation learning based on the experimental results in Section IV.
III-B Train with Simple and Single Object
@@
In this section, training strategy is outlined as: training the autoencoder depicted in Fig. 1 using a single simple object to significantly simplify data collection. However, the UniT representation obtained through this training method exhibits transferability and generalizability.
@@
In this paper, GelSight Mini with markers is used. We believe that other sensors from the “GelSight family” can also use our framework, as they share the same sensing principles. We chose the sensor pad with markers because it improves the tactile image’s capacity to capture the distribution of shear forces.
@@
The specific data collection process involves capturing images from the GelSight sensor at a fixed frequency (10 Hz), contacting the object with the GelSight, and continuously changing the contact configuration and the magnitude of the applied force.
@@
It is important to note that the reconstruction experiments conducted here are different from the depth reconstruction experiments shown in [7]. The goal of depth reconstruction is to generate depth images, whereas the purpose of image reconstruction here is to verify whether the learned representation contains sufficient information from the original image. Therefore, the results of depth reconstruction and image reconstruction are not comparable, since they are not the same type of experiments.
Our experimental results indicate that although UniT was trained only on a single simple object, the learned tactile representation can effectively generalize to unseen objects with diverse shapes, sizes, and textures. This tactile representation can reconstruct images that preserve most of the critical information of the original image, such as contact geometry and configuration (see Fig. 3). Moreover, the representation trained with an Allen key performs better than the one with the small ball. This is logical, as a small ball is one of the simplest objects from which to extract features: it lacks texture, edges, distinct shapes, and is omnidirectionally symmetrical. Despite this simplicity, the representation trained using a small ball can still capture features of unseen objects to a certain degree. Moreover, we have demonstrated that these representations can effectively generalize across different GelSight Minis, even though the training data originated from a single GelSight Mini.
@@
[…]
III-C Decoder Head for Downstream Tasks
After completion of the autoencoder training, both the decoder and the discriminator are discarded. The encoder is subsequently connected to the proposed decoder blocks and then to downstream tasks (such as perception and policy learning tasks).
As shown in Fig. 5, for decoding the representation $z\in\mathbb{R}^{h\times w\times c}$, we utilize a decoder block consisting of Conv2D, GroupNorm [37], SEBlock [38], and MaxPool2D as the basic module. This arrangement is designed to extract the information contained in the representation $z$.
In the experiments given in this paper, we adopt a strategy where, after the representation learning phase is completed, the pretrained encoder is frozen when applied to downstream tasks, and only the decoder blocks are trained. Although it is certainly possible to fine-tune the UniT pretrained encoder for specific downstream tasks, the results in Sections IV and V demonstrate that this approach of freezing the encoder for zero-shot transfer already yields sufficiently good performance. This underscores the robust generalizability and transferability of UniT, which are important when applying tactile representations on various perception and policy learning tasks with different objects (see Section V), as fine-tuning for each task and each object can be prohibitively costly. Robust generalizability and transferability are crucial to utilizing the tactile modality to achieve highly automated robotic agents.
IV Tactile Perception Experiments
BYOL MAE-ViT-Tiny MAE-ViT-Base UniT w/o VQ UniT
Mask Ratio Mask Ratio Representation Dimension Representation Dimension
0.25 0.5 0.75 0.25 0.5 0.75 $8\times 10$ $16\times 20$ $8\times 10$ $16\times 20$
Allen Key 1.33 0.337 0.358 0.374 0.247 0.273 0.370 0.225 0.189 0.156 0.128
Small Ball 1.06 0.435 0.685 0.789 0.526 0.560 0.388 0.283 0.185 0.274 0.166
TABLE I: USB plug 3D pose estimation results. Mean absolute error (in radians) on the test set of implementing representations learned from BYOL [19], MAE [18], and UniT. We test the representations trained on the Allen key dataset and the small ball dataset, respectively, as indicated in the top-right corner of Fig. 3. After the representation training phase, we freeze the encoders for all methods we benchmark here and train only the decoders on the 3D pose estimation dataset. The MAE with ViT-Base backbone has a model size of 143.20 M, whereas UniT with a representation dimension of $16\times 20$ has a model size of 79.81 M.
@@ IV Tactile Perception Experiments
TABLE II: USB plug 3D pose estimation results. Mean absolute error (in radians) on the test set of implementing ResNet [18], T3 [21], and UniT. For T3, we utilize the pretrained model released by the authors of [21], freezing the encoder in a manner similar to our approach with other representation learning methods. In the ResNet experiments, we train both the encoder and the decoders during training on 3D pose data.
@@
Figure 6: Overview of three tasks and hardware setup.
@@
TABLE III: Success rates of policy rollout on three different tasks. We use diffusion policy [2] as the policy backbone. For UniT, we use the encoder trained on the Allen Key dataset mentioned in Section III-B for GelSight Minis on both arms and for all three tasks. In addition, we also freeze the UniT encoder when integrating it into diffusion policy, as shown in Fig. 5. In (c), the heights on the left and right sides of the rack are 65/70 mm, 35/70 mm, and 50/70 mm. These varying heights result in different in-hand poses when the Allen key is grasped. The training dataset was collected using the 65/70 mm and 35/70 mm heights.
In this section, the effectiveness of UniT is demonstrated on a USB plug 3D pose estimation task and benchmark UniT with training a ResNet [18] from scratch, as well as other representation learning frameworks: BYOL [19], MAE [20], and the state-of-the-art tactile representation framework, T3 [21].
The task involves estimating the 3D pose of a USB plug based on its tactile image. As illustrated in Fig. LABEL:fig:first_page and our supplementary video, changes in the 3D pose of the USB plug result in altered contact configurations and corresponding variations in the tactile image. This type of 3D pose estimation is crucial for robotic insertion tasks that require high precision. This task is challenging because minor differences in the tactile images can correspond to significantly different 3D poses. The pose estimation model must accurately map these subtle tactile features to the correct 3D pose.
As shown in the supplementary video, we use the OptiTrack motion capture system to label the 3D pose groundtruth. In the data collection process, human move the USB plug randomly and we record the tactile raw image as well as 3D pose groundtruth in 10 Hz.

### v2

# UniT: Data Efficient Tactile Representation with Generalization to Unseen Objects

@@ Abstract
UniT is an approach to tactile representation learning, using VQGAN to learn a compact latent space and serve as the tactile representation. It uses tactile images obtained from a single simple object to train the representation with generalizability. This tactile representation can be zero-shot transferred to various downstream tasks, including perception tasks and manipulation policy learning. Our benchmarkings on in-hand 3D pose and 6D pose estimation tasks and a tactile classification task show that UniT outperforms existing visual and tactile representation learning methods. Additionally, UniT’s effectiveness in policy learning is demonstrated across three real-world tasks involving diverse manipulated objects and complex robot-object-environment interactions. Through extensive experimentation, UniT is shown to be a simple-to-train, plug-and-play, yet widely effective method for tactile representation learning. For more details, please refer to our open-source repository https://github.com/ZhengtongXu/UniT and the project website https://zhengtongxu.github.io/unit-website/.
@@ I Introduction
Imitation learning has shown promising results equipping robots in the domain of manipulation, with the ability to master complex, high-precision, dexterous skills through human demonstrations [1, 2]. However, the predominant focus of current research is on leveraging image or point cloud inputs to perceive scene information. During manipulation, robots encounter varied force interactions from both the environments and the manipulated objects. The sole dependence on images [1, 2] or point clouds [3] could obscure critical details about the objects’ in-hand states and the dynamics of force interactions, which may be crucial for effective manipulation. Despite the current focus on visual information in robot learning research, humans routinely use visual and tactile feedback in manipulation tasks. Therefore, exploring the integration of visual and tactile modalities in imitation learning could potentially enhance the robots’ performance in manipulation tasks.
Vision-based tactile sensors like GelSight [4] provide high-resolution feedback, capturing rich tactile information. Prior research suggests that GelSight’s unique sensing principles enable trained neural networks to generalize effectively. For instance, a gradient estimation model trained solely on images of a small ball contacting the sensor generalizes across objects with various shapes and textures [5]. Likewise, a tactile-reactive grasping policy trained on standardized texture-less blocks extends to diverse objects [6]. While these studies highlight the distinctiveness of tactile images compared to standard visual images, their generalizability remains task-specific. This naturally raises the following question: can we use a single simple object to learn a data efficient tactile representation of one type of GelSight that 1) possesses generalizability, 2) incorporates as much of the rich information present in tactile images as possible, and 3) can be applied in a zero-shot manner across a variety of downstream tasks involving different objects?
To address this question, we propose UniT, data efficient tactile representation with generalization to unseen objects. Our contributions are summarized as follows:
@@
1. Easy to Train yet Broadly Applicable: We introduce UniT, a tactile representation learning method for GelSight sensors. UniT uses VQGAN to learn a structured latent space and serve as the tactile representation. It can utilize data on a single simple object to learn a data efficient tactile representation that can be generalized to objects of different sizes and shapes. Through experiments in image reconstruction involving diverse objects, we demonstrate that the representation learned by UniT from a single simple object can capture information of unseen objects on contact configurations, object shapes, and dynamic marker motions induced by applied forces. This kind of generalizability makes UniT both straightforward to train and broadly applicable.
@@
2. Deploy for Tactile Perception: The tactile encoder trained through UniT can be seamlessly transferred to downstream tactile perception tasks. Furthermore, due to UniT’s lenient data requirements during training, representations learned even from the simplest objects, such as a small ball, can effectively facilitate tactile perception of everyday objects. In experiments involving the task of estimating the in-hand 6D pose of the USB plug, UniT outperforms training of a ResNet [7] from scratch, existing visual representation learning methods BYOL [8] and MAE [9], and the state-of-the-art tactile representation learning method, T3 [10].
@@ II-A Tactile-involved Imitation Learning
Research that incorporates tactile arrays [11, 12] and force-torque feedback [13] into imitation learning policies have shown strong performance in tasks with complex interactions between environments, objects, and robots. However, tactile interactions are inherently complex and these modalities often do not capture the full spectrum of tactile information. For example, tactile arrays may struggle to accurately perceive an object’s geometry, in-hand position, and orientation, while force-torque may not effectively capture dynamic force distributions on the hand/finger.
Additionally, vision-based tactile sensors such as GelSight [4] offer high-resolution feedback that captures extensive tactile information. Recent studies incorporating GelSight into imitation learning have demonstrated its considerable potential and enhanced performance [14, 15].
@@ II-B Representation Learning in Imitation Learning and Tactile Sensing
Recent studies demonstrate that representation learning significantly enhances performance in imitation learning [16, 17]. Encoders that are pretrained with representation learning are particularly adept at extracting information from visual observations.
In the field of tactile sensing, a variety of representation learning frameworks have been introduced, including the use of MAE [18, 19] and CNN [20]. Furthermore, the works in [10, 21] introduce frameworks for tactile representation learning that scales across multi-sensors and multi-tasks. The work in [22] learns multimodal tactile representation by aligning tactile embeddings to pretrained image embeddings associated with a variety of other modalities.
In this paper, UniT demonstrates that tactile images from a single, simple object can train a highly generalizable tactile representation—a capability absent in existing work.
III Background: VQGAN autoencoder
The work in [23] introduces VQGAN for high-resolution image synthesis. VQGAN integrates a VQVAE [24] with a patch-based discriminator. Subsequently, the work in [25] introduces latent diffusion model, which leverages the VQGAN autoencoder to compress images into a latent space where diffusion and denoising processes can occur. The efficacy of the latent diffusion model is largely due to the robust image compression capabilities of VQGAN, which not only facilitates high-quality image reconstruction but also avoids high degree of variance in the learned latent space through VQ regularization.
Although VQGAN demonstrates impressive performance in generative model and image generation areas, its role as an autoencoder is primarily to enable high-quality image compression and reconstruction. In this paper, we demonstrate that VQGAN can serve as a highly effective tactile representation learner that can be trained with minimal data.
IV Method
[…]
@@ IV Method
[…]
Figure 6: Visualization of latent spaces. We transformed tactile images of size 3×128×160 into latent spaces of size 3×16×20 and visualized these latent spaces as RGB images. Both VQGAN and CNN autoencoders are trained solely on the Allen key dataset. We tested three different unseen objects. The three images on the left are from artificial strawberry, the three in the middle are from nut, and the three on the right are from screw.
IV-A Training Pipeline
Tactile images are created based on a unique imaging principle that combines RGB-tricolor light scattering with gel deformation. Unlike visual images, however, tactile images exhibit a much more compact color distribution. This compactness arises because the imaging process disregards object and background colors, focusing solely on tactile features such as contact configuration, geometry, and force distribution. Given the inherently low-variance color distribution in raw tactile images, we believe that VQGAN can effectively learn a mapping from raw tactile images to a latent space, and that this mapping can be achieved with a small amount of tactile data. Moreover, since VQGAN is designed to learn a low-variance latent space, we propose that for tactile images, such a latent space is not only sufficient to represent the raw tactile images but also enhances data efficiency and applicability for downstream tasks due to its structured characteristics. Accordingly, we leverage VQGAN for learning tactile representations.
@@ IV-A Training Pipeline
As depicted in Fig. 1, the representation learning pipeline employs a VQGAN architecture with the quantization layer absorbed into the decoder. This pipeline comprises a CNN encoder ${E}$, a CNN decoder incorporating the quantization layer ${G}$, and a patch-based discriminator ${D}$. The entire framework is trained in a self-supervised manner. Denote the tactile image input as $x\in\mathbb{R}^{{H}\times{W}\times{3}}$, where the encoder ${E}$ maps the image to a tactile representation $z={E}(x)\in\mathbb{R}^{{h}\times{w}\times{c}}$. The decoder reconstructs $z$ to an image $\hat{x}$.
@@
The use of a discriminator is justified because recent studies [26] have shown that adversarial loss can enhance the performance of representation learning. The presence of VQ regularization enables the learning of a structured tactile representation with reduced variance, which can improve the performance of representation learning based on the experimental results in Section VI.
IV-B Train with Simple and Single Object
@@
In this section, training strategy is outlined as: training the autoencoder depicted in Fig. 1 using a single simple object to significantly simplify data collection. However, the UniT representation obtained through this training method exhibits generalizability.
@@
In this paper, GelSight mini with markers is used. We believe that other sensors from the “GelSight family” can also use our framework, as they share the same sensing principles. We chose the sensor gel with markers because it improves the tactile image’s capacity to capture the distribution of shear forces.
@@
The specific data collection process involves capturing images from the GelSight sensor at a fixed frequency 10 Hz, contacting the object with the GelSight, and continuously changing the contact configuration and the magnitude of the applied force.
@@
IV-C Decoder Head for Downstream Tasks
After completion of the autoencoder training, both the decoder and the discriminator are discarded. The encoder is subsequently connected to the proposed decoder blocks and then to downstream tasks like perception and policy learning tasks.
As shown in Fig. 2, for decoding the representation $z\in\mathbb{R}^{h\times w\times c}$, we utilize a decoder block consisting of Conv2D, GroupNorm [27], SEBlock [28] as the basic module.
In the experimental section of this paper, we demonstrate that after the representation learning phase is complete, freezing the UniT encoder and training only the decoder head for downstream tasks achieves performance comparable to fine-tuning the entire model.
V Reconstruction Experiments
@@
Our experimental results indicate that although UniT was trained only on a single simple object, the learned tactile representation can effectively generalize to unseen objects with diverse shapes, sizes, and textures. This tactile representation can reconstruct images that preserve most of the critical information of the original image, such as contact geometry and configuration, as shown in Fig. 4. Moreover, the representation trained with an Allen key performs better than the one with the small ball. This is logical, as a small ball is one of the simplest objects from which to extract features: it lacks texture, edges, distinct shapes, and is omnidirectionally symmetrical. Despite this simplicity, the representation trained using a small ball can still capture features of unseen objects to a certain degree.
As shown in Fig. 4, the representation learned by UniT can generalize across different units of the same type of sensor for image reconstruction, even when the representation is trained on a single object with one sensor unit. UniT does not support transfer across different sensor types with drastically different image characteristics, as it relies on minimal data from a single sensor type. However, UniT’s data collection and training are efficient, requiring only a small dataset from one object on a single sensor. Once trained, the representation can be applied to diverse downstream tasks of this type of sensor. More results of downstream tasks will be shown in Sections VI and VII.
@@
We also conduct an experiment to visualize the latent spaces, providing an illustration of the working mechanism of our proposed tactile representation, as shown in Fig. 6. Specifically, we train a pure CNN-based autoencoder and a VQGAN autoencoder on Allen key data and use their respective encoders to project the same original image into different latent spaces. Specifically, we transformed tactile images of size 3×128×160 into latent spaces of size 3×16×20 and visualized these latent spaces as RGB images. The results reveal key differences: without VQ, the latent space retains a color distribution highly similar to the original image. This outcome is expected, as reconstruction training without VQ relies on a pure CNN-based autoencoder, where convolutions inherently preserve the original color distribution. In contrast, the VQGAN-based latent space exhibits a significantly more structured representation. While preserving essential characteristics of the original image, VQGAN amplifies meaningful features while suppressing background color variations, resulting in a more distinct, dichotomous visualization. Unlike the CNN-based latent space, which retains background color variance, the VQGAN-encoded representation focuses primarily on salient features, aligning with its ability to learn a low-variance latent representation.
VI Tactile Perception Experiments
BYOL MAE-ViT-Tiny MAE-ViT-Base UniT w/o VQ UniT w/o Dis. UniT
Mask Ratio Mask Ratio Rep. Dim. Rep. Dim. Rep. Dim.
0.25 0.5 0.75 0.25 0.5 0.75 $8\times 10$ $16\times 20$ $8\times 10$ $16\times 20$ $8\times 10$ $16\times 20$
A. K. 1.33 0.337 0.358 0.374 0.247 0.273 0.370 0.225 0.189 0.190 0.145 0.156 0.128
S. B. 1.06 0.435 0.685 0.789 0.526 0.560 0.388 0.283 0.185 0.261 0.193 0.274 0.166
@@ VI Tactile Perception Experiments
TABLE I: USB plug 3D pose estimation results. The metric reported here is the mean absolute error across the entire test set, with the unit in radians.
BYOL MAE T3 ResNet UniT
0.25 0.5 0.75 Tiny Small Medium Large Pretrain Random
Rotation 1.202 0.310 0.235 0.284 1.551 0.914 0.306 0.338 0.336 0.365 0.155
Position 11.2 6.2 6.5 6.0 11.7 10.8 5.8 5.7 5.8 6.6 4.8
TABLE II: USB plug 6D pose estimation results. We report the mean absolute error for both rotation (radians) and position (millimeters) across the entire test set.
BYOL MAE T3 ResNet UniT
0.25 0.5 0.75 Tiny Small Medium Large Pretrain Random
Freeze (%) 85.0 77.1 77.1 81.0 0 35.4 82.3 79.2 85.4 72.1 92.1
Fine-Tune (%) 97.5 85.4 81.3 89.6 89.6 87.7 85.4 85.0 97.9 85.4 97.3
TABLE III: Classification benchmarks on YCB-Sight dataset [29]. The representations for BYOL, MAE, and UniT are all obtained using training data from all six objects.
all master chef can sugar box tomato soup can potted meat can bleach cleanser wood block
Freeze (%) 92.1 87.9 90.0 86.3 87.9 91.0 89.0
Fine-Tune(%) 97.3 90.4 88.5 87.7 87.7 93.5 92.5
TABLE IV: UniT representation training ablation study. “all” indicates the representation trained on all six objects, while other columns, such as “sugar box,” refer to representations trained only on data from that specific object.
@@
Figure 7: Overview of three tasks and hardware setup. In (c), the heights on the left and right sides of the rack are 65/70 mm, 35/70 mm, and 50/70 mm. These varying heights result in different in-hand poses when the Allen key is grasped. The training dataset was collected using the 65/70 mm and 35/70 mm heights.
@@
TABLE V: Success rates of policy rollout on three different real tasks.
Vision Only Visual Tactile from Scratch T3 Freeze T3 Fine-Tune UniT Freeze UniT Fine-Tune
Success Rate 0.501 0.502 0.509 0.516 0.571 0.575
TABLE VI: Success rate of policy rollout on the simulated peg insertion task shown in Fig. 8.
[…]
Figure 8: Simulated peg insertion task using TacSL [30].
In this section, the effectiveness of UniT is demonstrated on USB plug 3D pose and 6D pose estimation tasks, and tactile classification task. We benchmark UniT with training a ResNet [7] from scratch, as well as other representation learning frameworks: BYOL [8], MAE [9], and the state-of-the-art tactile representation framework, T3 [10]. In all experiments, we evaluate performance on the test set every 10 training epochs and report the average metric values from the last ten checkpoints.
VI-A Pose Estimation
The pose estimation tasks involve estimating the 3D or 6D pose of a USB plug based on its tactile image. Here, we conduct two types of experiments: in the 3D pose task, we estimate only the rotation, and the model output is a 4-dimensional quaternion. In the 6D pose task, we estimate both position and rotation simultaneously, with the model outputting a 7-dimensional vector (3-dimensional position + 4-dimensional quaternion). As illustrated in our supplementary video, changes in the 6D pose of the USB plug result in altered contact configurations and corresponding variations in the tactile image. This type of pose estimation is crucial for robotic insertion tasks that require high precision. This task is challenging because minor differences in the tactile images can correspond to significantly different poses. The pose estimation model must accurately map these subtle tactile features to the correct pose.
As shown in the supplementary video, we use the OptiTrack motion capture system to label the 6D pose ground truth. In the data collection process, human move the USB plug randomly and we record the raw tactile image as well as 6D pose ground truth in 10 Hz.